# EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN)

### v1

@@
[Uncaptioned image] CERN-LHCb-DP-2024-003 March 6, 2025
@@ 1 Introduction
In the RICH1 detector, the MaPMTs are operated by using a powering scheme involving the bias of the last of the twelve dynodes, in order to maintain a linear response of the MaPMTs with the expected rates, as described in Sec. 2. The implementation of the method is described in Sec. 3, where it is shown how an output signal proportional to the luminosity is obtained through the measurement of the power-supply currents. The possibility to use the number of Cherenkov hits for the same purpose is verified through simulated samples generated with different pile-up conditions, as described in Sec. 4, and calibrated through van der Meer scans performed during the 2023 data-taking period as shown in Sec. 5. In order to calibrate the MaPMT anode currents to provide a real-time luminosity signal, a data-taking phase during which the instantaneous luminosity requested to the LHC is changed in known steps is used and combined with the information coming from the number of Cherenkov hits, described in Sec. 6. Finally, the results and prospects are discussed in Sec. 7.
@@ 2 Modelling of the MaPMT response
For tube HPK-R11265-M64, equipped with 12 dynodes, the standard voltage divider suggested by the manufacturer, shown in Fig. 2, is non-tapered, except for the first (from photocathode, labelled PK) and last stage, and is optimised for maximum gain for single-photon detection:
$(\text{PK})\qquad R_{0}\{2.3,1.2,1,1,1,1,1,1,1,1,1,1,1.5\}\qquad$ (1)
Using $R_{0}=200~{}\mathrm{k\Omega}$, the total resistance of the voltage divider is $R=3000~{}\mathrm{k\Omega}$. The anode is grounded and a constant negative voltage, typically $V_{\mathrm{1}}=-900~{}\mathrm{V}$, is applied to the cathode resulting in an average gain of order $G=10^{6}$ electrons ($1~{}\mathrm{Me}$) and, in absence of a cathode photocurrent, in a power-supply current $i_{\text{PS}}=300~{}\mathrm{\mu A}$ flowing through the resistors of the voltage divider. [^1] Similar considerations to those made in this section can be done for tube HPK-R12699-M64, equipped with 10 dynodes, having a voltage divider $R_{0}\{2,1,1,1,1,1,1,1,1,1,1.5\}$ for a total resistance of $R=2500~{}\mathrm{k\Omega}$.

### v2

@@
[Uncaptioned image] CERN-LHCb-DP-2024-003 July 14, 2025
@@ 1 Introduction
In the RICH1 detector, the MaPMTs are operated by using a powering scheme involving the bias of the last of the twelve dynodes, in order to maintain a linear response of the MaPMTs with the expected rates, as described in Sec. 2. The implementation of the method is described in Sec. 3, where it is shown how an output signal proportional to the luminosity is obtained through the measurement of the power-supply currents. The possibility to use the number of Cherenkov hits for the same purpose is verified through simulated samples generated with different pile-up conditions, as described in Sec. 4, and calibrated through van der Meer scans performed during the 2023 data-taking period as shown in Sec. 5. To calibrate the MaPMT anode currents for providing a real-time luminosity signal, a dedicated data-taking phase is performed in which the instantaneous luminosity requested from the LHC is varied in known steps, as described in Sec. 6. Finally, the results and future prospects are presented in Sec. 7.
@@ 2 Modelling of the MaPMT response
For tube HPK-R11265-M64, equipped with 12 dynodes, the standard voltage divider suggested by the manufacturer, shown in Fig. 2, is non-tapered, except for the first (from photocathode, labelled PK) and last stage, and is optimised for maximum gain for single-photon detection. The total resistance of the voltage divider is $R=3000~{}\mathrm{k\Omega}$.
The anode is grounded and a constant negative voltage, typically $V_{\mathrm{1}}=-900~{}\mathrm{V}$, is applied to the cathode resulting in an average gain of order $G=10^{6}$ electrons ($1~{}\mathrm{Me}$) and, in absence of a cathode photocurrent, in a power-supply current $i_{\text{PS}}=300~{}\mathrm{\mu A}$ flowing through the resistors of the voltage divider. [^1] Similar considerations to those made in this section can be done for tube HPK-R12699-M64, equipped with 10 dynodes and having a total resistance of $R=2500~{}\mathrm{k\Omega}$.